OpenStudy (anonymous):

Susie has designed an exercise program for herself. One part of the program requires her to walk between 25 and 30 miles each week. She plans to walk the same distance each day five days a week. What is the range of miles that she should walk each day?

OpenStudy (anonymous):

I got 5-6 miles, but I don't know if that's right.

OpenStudy (anonymous):

i come here for the metal. Lowest :25 Highest:35 each day x miles keep x=25/5 x=35/5 [5,7]

OpenStudy (anonymous):

sorry i revised